OpenStudy (anonymous):

Find the distance between these points. A (5, 8), B(-3, 4) AB =

OpenStudy (anonymous):

A(5, 8), B(-3, 4)Distance AB = square root of [(x2-x1)^2 +(y2-y1)^2] = sq. root of [(-3-5)^2+(4-8)^2] = square root of 80 or 8.944 ...

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)